# Asymptotics for conformal inference

### v1

@@ Appendix B Proofs of auxiliary results
$\begin{pmatrix}\sqrt{m_{0}(m)}\left[{\widehat{F}}_{{\tiny\mbox{$m$,$0$,test}}}-I\right]\\ \sqrt{m_{1}(m)}\left[{\widehat{F}}_{{\tiny\mbox{$m$,$1$,test}}}-{F_{\tiny\mbox{test,cal}}}\right]\\ \sqrt{n}\left[{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-(1-I)\right]\\ \sqrt{n}\left[{F_{\tiny\mbox{test,cal}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-{F_{\tiny\mbox{test,cal}}}(1-I)\right]\\ {({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\\ \mathbb{V}(1-{F_{\tiny\mbox{test,cal}}})\\ \mathbb{W}\\ G^{\prime}\mathbb{W}\\ 1-I\end{pmatrix}\text{ on $\left[D(0,1)\right]^{4}\times\ell^{\infty}(0,1)$},$
where we denoted ${\widehat{F}}_{{\tiny\mbox{$m$,$r$,test}}}$ the e.c.d.f. of $\left\{T_{i},i\in\llbracket m\rrbracket\cap{\mathcal{H}}_{r}\right\}$ for $r\in\left\{0,1\right\}$. Now, by using Lemma E.10 (or more precisely, an obvious extension of it for $5$ joint processes), we obtain
$\begin{pmatrix}\sqrt{m_{0}(m)}\left[{\widehat{F}}_{{\tiny\mbox{$m$,$0$,test}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)\right]\\ \sqrt{m_{1}(m)}\left[{\widehat{F}}_{{\tiny\mbox{$m$,$1$,test}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-{F_{\tiny\mbox{test,cal}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)\right]\\ \sqrt{n}\left[{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-(1-I)\right]\\ \sqrt{n}\left[{F_{\tiny\mbox{test,cal}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-{F_{\tiny\mbox{test,cal}}}(1-I)\right]\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\\ \mathbb{V}\circ G\\ \mathbb{W}\\ G^{\prime}\mathbb{W}\end{pmatrix}\text{ on $\left[D(0,1)\right]^{4}$}.$
@@
$\sqrt{\tau_{n,m}}\left[{\widehat{{G}}}_{m,{0}}^{(n)}-I\right]=$ $-{\sqrt{\frac{\tau_{n,m}}{m}\times\frac{m}{m_{0}}}}\sqrt{m_{0}}\left[{\widehat{F}}_{{\tiny\mbox{$m$,$0$,test}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)\right]$
$-\sqrt{\frac{\tau_{n,m}}{n}}\sqrt{n}\left[{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-(1-I))\right],$
$\sqrt{\tau_{n,m}}\left[{\widehat{{G}}}_{m,{1}}^{(n)}-G\right]=$ $-{\sqrt{\frac{\tau_{n,m}}{m}\times\frac{m}{m_{1}}}}\sqrt{m_{1}}\left[{\widehat{F}}_{{\tiny\mbox{$m$,$1$,test}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-{F_{\tiny\mbox{test,cal}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)\right]$
$-\sqrt{\frac{\tau_{n,m}}{n}}\sqrt{n}\left[{F_{\tiny\mbox{test,cal}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-{F_{\tiny\mbox{test,cal}}}(1-I)\right].$
@@
By Assumption 5, and since $\alpha>[G_{{\tiny\mbox{mixt}}}^{\prime}(0^{+})]^{-1}$, the asymptotic threshold $\mathcal{T}_{\alpha}=\mathcal{T}^{\mathrm{BH}_{\alpha}}(G_{{\tiny\mbox{mixt}}})$ is well defined and belongs to $(0,1)$. Now observe that there exists a compact interval $[a,b]\subset(0,1)$ such that
$\mathcal{T}_{\alpha}=\sup\left\{t\in[a,b],\ G_{{\tiny\mbox{mixt}}}(t)\geqslant\frac{t}{\alpha}\right\},$ (44)
@@
$\mathcal{T}^{\mathrm{BH}_{\alpha}}_{[a,b]}(F)\coloneqq\sup\left\{t\in[a,b],F(t)\geqslant\frac{t}{\alpha}\right\},$ (45)
@@
$\left\{\left|{\widehat{G}}_{m}^{(n)}(a)-G_{{\tiny\mbox{mixt}}}(a)\right|\leqslant\eta\right\}\subset\left\{\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{(n)})=\mathcal{T}^{\mathrm{BH}_{\alpha}}_{[a,b]}({\widehat{G}}_{m}^{(n)})\right\},$ (46)
@@
$\sqrt{\tau_{n,m}}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}_{[a,b]}({\widehat{G}}_{m}^{(n)})-\mathcal{T}_{\alpha}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{1}{\frac{1}{\alpha}-G_{{\tiny\mbox{mixt}}}^{\prime}(\mathcal{T}_{\alpha})}\mathbb{Z}({\mathcal{T}_{\alpha}});$
$\sqrt{\tau_{n,m}}\left(\mathrm{FDP}_{m}^{(n)}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}_{[a,b]}({\widehat{G}}_{m}^{(n)})\right)-\pi_{0}\alpha\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{\pi_{0}}{G_{{\tiny\mbox{mixt}}}(\mathcal{T}_{\alpha})}\mathbb{Z}_{0}({\mathcal{T}_{\alpha}});$
$\sqrt{\tau_{n,m}}\left(\mathrm{TDP}_{m}^{(n)}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}_{[a,b]}({\widehat{G}}_{m}^{(n)})\right)-G(\mathcal{T}_{\alpha})\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{G^{\prime}(\mathcal{T}_{\alpha})}{\frac{1}{\alpha}-G_{{\tiny\mbox{mixt}}}^{\prime}(\mathcal{T}_{\alpha})}\mathbb{Z}({\mathcal{T}_{\alpha}})+\mathbb{Z}_{1}(\mathcal{T}_{\alpha}),$
@@ Appendix C Proofs for the weighted case
${({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}\overset{\mathbb{P}}{\rightarrow}({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}\text{ on $\ell^{\infty}(0,1)$,}$
@@
$\sqrt{n}\left({({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}-({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}\right)\overset{\mathcal{L}}{\rightarrow}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1})^{\prime}{\rho^{w}}\left(\mathbb{V}\left({{V}^{w}_{{\tiny\mbox{cal}}}}\circ({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}\right)+\left[I-{{V}^{w}_{{\tiny\mbox{cal}}}}\circ({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}\right]N\right)$
[…]
@@
Since $w$ is uniformly bounded, the family =$\mathcal{F}=\left\{w_{t}:x\in\mathbb{R}\mapsto w(x){\mathds{1}_{x\leqslant t}};t\in\mathbb{R}\right\}$ is $\mathcal{U}(0,1)$-Glivenko-Cantelli, and $(K^{(n)})_{n}$ converges uniformly on $[0,1]$ in probability to the function $K$ given by $K(t)\coloneqq\int_{0}^{t}w(x)\operatorname{d}x$ (Shorack and Wellner,, 1986). Then, by continuity at $K$ of the map of Lemma E.8, we get that $(K^{(n)}/K^{(n)}(1))_{n}$ converges uniformly (in probability) to ${F^{w}_{{\tiny\mbox{cal}}}}$. Since $\|{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}-K^{(n)}/K^{(n)}(1)\|_{\infty}\leqslant w(+\infty)/((n+1)K^{(n)}(1)+w(+\infty))$ tends uniformly to 00 a.s., $({\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)})_{n}$ converges uniformly (in probability) to ${F^{w}_{{\tiny\mbox{cal}}}}$. By continuity of the inverse map at ${F^{w}_{{\tiny\mbox{cal}}}}$ (see Lemma E.6), we obtain that $({({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}})_{n}$ converges in probability to $({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}$ on $\ell^{\infty}(0,1)$. This proves the first statement.
Next, we turn to prove the second statement. Since $w$ is uniformly bounded, the family =$\mathcal{F}=\left\{w_{t}:x\in\mathbb{R}\mapsto w(x){\mathds{1}_{x\leqslant t}};t\in\mathbb{R}\right\}$ is $\mathcal{U}(0,1)$-Donsker (Shorack and Wellner,, 1986). hence there exist $\mathbb{K}=(\mathbb{K}(t))_{t\in[0,1]}$ a Gaussian process such that
$\sqrt{n}\left(K^{(n)}-K\right)\overset{\mathcal{L}}{\rightarrow}\mathbb{K}\ \text{in $\ell^{\infty}[0,1]$},$
@@
$\mbox{Cov}\left(\mathbb{K}(s),\mathbb{K}(t)\right)$ $=\int_{0}^{s\wedge t}w^{2}(x)\operatorname{d}x-\int_{0}^{s}w(x)\operatorname{d}x\int_{0}^{t}w(x)\operatorname{d}x;$
$=K(1)^{2}\left({\rho^{w}}^{2}{{{V}^{w}_{{\tiny\mbox{cal}}}}}(s\wedge t)-{{F^{w}_{{\tiny\mbox{cal}}}}}(t){{F^{w}_{{\tiny\mbox{cal}}}}}(s)\right),$
@@
$\sqrt{n}\left(\left(\frac{K^{(n)}}{K^{(n)}(1)}\right)-{{F^{w}_{{\tiny\mbox{cal}}}}}\right)\overset{\mathcal{L}}{\rightarrow}\frac{\mathbb{K}}{K(1)}-\frac{\mathbb{K}(1)}{K(1)}{{F^{w}_{{\tiny\mbox{cal}}}}},\ \text{in $\mathcal{C}[0,1]$}.$
@@
$\sqrt{n}\left({\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}-{{F^{w}_{{\tiny\mbox{cal}}}}}\right)\overset{\mathcal{L}}{\rightarrow}\frac{\mathbb{K}}{K(1)}-\frac{\mathbb{K}(1)}{K(1)}{{F^{w}_{{\tiny\mbox{cal}}}}},\ \text{in $\mathcal{C}[0,1]$}.$
@@
$\sqrt{n}\left({\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)-1}-({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}\right)\overset{\mathcal{L}}{\rightarrow}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1})^{\prime}\left(-\frac{\mathbb{K}\circ({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}}{K(1)}+\frac{\mathbb{K}(1)}{K(1)}I\right),\ \text{in $D(0,1)$}.$
To conclude, we identify the covariance of the Gaussian limiting process. For, $01>t\geqslant s>$ 0,
$\mbox{Cov}\left(-\mathbb{K}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(t))/K(1)+\mathbb{K}(1)/K(1)t,-\mathbb{K}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(s))/K(1)+\mathbb{K}(1)/K(1)s\right)$
@@
$={\rho^{w}}^{2}\left({{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(s))-s{{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(t))-t{{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(s))+ts\right).$
@@
$+\left[t-{{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(t))\right]\left[s-{{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(t))\right]\Big{)}$
$=$ ${\rho^{w}}^{2}\left({{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(s))\left[1-{{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(t))\right]+\left[t-{{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(t))\right]\left[s-{{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(t))\right]\right).$
[…]

### v3

@@ Appendix B Proofs of auxiliary results
$\begin{pmatrix}\sqrt{m_{0}(m)}\mathopen{}\mathclose{{\left[{\widehat{F}}_{{\tiny\mbox{$m$,$0$,test}}}-I}}\right]\\ \sqrt{m_{1}(m)}\mathopen{}\mathclose{{\left[{\widehat{F}}_{{\tiny\mbox{$m$,$1$,test}}}-{F_{\tiny\mbox{test,cal}}}}}\right]\\ \sqrt{n}\mathopen{}\mathclose{{\left[{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-(1-I)}}\right]\\ \sqrt{n}\mathopen{}\mathclose{{\left[{F_{\tiny\mbox{test,cal}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-{F_{\tiny\mbox{test,cal}}}(1-I)}}\right]\\ {({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\\ \mathbb{V}(1-{F_{\tiny\mbox{test,cal}}})\\ \mathbb{W}\\ G^{\prime}\mathbb{W}\\ 1-I\end{pmatrix}\text{ on $\mathopen{}\mathclose{{\left[D(0,1)}}\right]^{4}\times\ell^{\infty}(0,1)$},$
where we denoted ${\widehat{F}}_{{\tiny\mbox{$m$,$r$,test}}}$ the e.c.d.f. of $\mathopen{}\mathclose{{\left\{T_{i},i\in\llbracket m\rrbracket\cap{\mathcal{H}}_{r}}}\right\}$ for $r\in\mathopen{}\mathclose{{\left\{0,1}}\right\}$. Now, by using Lemma E.10 (or more precisely, an obvious extension of it for $5$ joint processes), we obtain
$\begin{pmatrix}\sqrt{m_{0}(m)}\mathopen{}\mathclose{{\left[{\widehat{F}}_{{\tiny\mbox{$m$,$0$,test}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)}}\right]\\ \sqrt{m_{1}(m)}\mathopen{}\mathclose{{\left[{\widehat{F}}_{{\tiny\mbox{$m$,$1$,test}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-{F_{\tiny\mbox{test,cal}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)}}\right]\\ \sqrt{n}\mathopen{}\mathclose{{\left[{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-(1-I)}}\right]\\ \sqrt{n}\mathopen{}\mathclose{{\left[{F_{\tiny\mbox{test,cal}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-{F_{\tiny\mbox{test,cal}}}(1-I)}}\right]\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\\ \mathbb{V}\circ G\\ \mathbb{W}\\ G^{\prime}\mathbb{W}\end{pmatrix}\text{ on $\mathopen{}\mathclose{{\left[D(0,1)}}\right]^{4}$}.$
@@
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left[{\widehat{{G}}}_{m,{0}}^{(n)}-I}}\right]=$ $-{\sqrt{\frac{\tau_{n,m}}{m}\times\frac{m}{m_{0}}}}\sqrt{m_{0}}\mathopen{}\mathclose{{\left[{\widehat{F}}_{{\tiny\mbox{$m$,$0$,test}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)}}\right]$
$-\sqrt{\frac{\tau_{n,m}}{n}}\sqrt{n}\mathopen{}\mathclose{{\left[{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-(1-I))}}\right],$
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left[{\widehat{{G}}}_{m,{1}}^{(n)}-G}}\right]=$ $-{\sqrt{\frac{\tau_{n,m}}{m}\times\frac{m}{m_{1}}}}\sqrt{m_{1}}\mathopen{}\mathclose{{\left[{\widehat{F}}_{{\tiny\mbox{$m$,$1$,test}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-{F_{\tiny\mbox{test,cal}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)}}\right]$
$-\sqrt{\frac{\tau_{n,m}}{n}}\sqrt{n}\mathopen{}\mathclose{{\left[{F_{\tiny\mbox{test,cal}}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-{F_{\tiny\mbox{test,cal}}}(1-I)}}\right].$
@@
By Assumption 5, and since $\alpha>[G_{{\tiny\mbox{mixt}}}^{\prime}(0^{+})]^{-1}$, the asymptotic threshold $\mathcal{T}_{\alpha}=\mathcal{T}^{\mathrm{BH}_{\alpha}}(G_{{\tiny\mbox{mixt}}})$ is well defined and belongs to $(0,1)$. Now observe that there exists a compact interval $a,b]\subset(0,1)$ such that
$\mathcal{T}_{\alpha}=\sup\mathopen{}\mathclose{{\left\{t\in[a,b],\ G_{{\tiny\mbox{mixt}}}(t)\geqslant\frac{t}{\alpha}}}\right\},$ (50)
@@
$\mathcal{T}^{\mathrm{BH}_{\alpha}}_{[a,b]}(F)\coloneqq\sup\mathopen{}\mathclose{{\left\{t\in[a,b],F(t)\geqslant\frac{t}{\alpha}}}\right\},$ (51)
@@
$\mathopen{}\mathclose{{\left\{\mathopen{}\mathclose{{\left|{\widehat{G}}_{m}^{(n)}(a)-G_{{\tiny\mbox{mixt}}}(a)}}\right\rvert\leqslant\eta}}\right\}\subset\mathopen{}\mathclose{{\left\{\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{(n)})=\mathcal{T}^{\mathrm{BH}_{\alpha}}_{[a,b]}({\widehat{G}}_{m}^{(n)})}}\right\},$ (52)
@@
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}_{[a,b]}({\widehat{G}}_{m}^{(n)})-\mathcal{T}_{\alpha}}}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{1}{\frac{1}{\alpha}-G_{{\tiny\mbox{mixt}}}^{\prime}(\mathcal{T}_{\alpha})}\mathbb{Z}({\mathcal{T}_{\alpha}});$
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{FDP}_{m}^{(n)}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}_{[a,b]}({\widehat{G}}_{m}^{(n)})}}\right)-\pi_{0}\alpha}}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{\pi_{0}}{G_{{\tiny\mbox{mixt}}}(\mathcal{T}_{\alpha})}\mathbb{Z}_{0}({\mathcal{T}_{\alpha}});$
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{TDP}_{m}^{(n)}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}_{[a,b]}({\widehat{G}}_{m}^{(n)})}}\right)-G(\mathcal{T}_{\alpha})}}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{G^{\prime}(\mathcal{T}_{\alpha})}{\frac{1}{\alpha}-G_{{\tiny\mbox{mixt}}}^{\prime}(\mathcal{T}_{\alpha})}\mathbb{Z}({\mathcal{T}_{\alpha}})+\mathbb{Z}_{1}(\mathcal{T}_{\alpha}),$
@@ Appendix C Proofs for the weighted case
[…]
@@
$\sqrt{n}\mathopen{}\mathclose{{\left({({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}-({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}}}\right)\overset{\mathcal{L}}{\rightarrow}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1})^{\prime}{\rho^{w}}\mathopen{}\mathclose{{\left(\mathbb{V}\circ{{{V}^{w}_{{\tiny\mbox{cal}}}}\circ({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}}+\mathopen{}\mathclose{{\left[I-{{V}^{w}_{{\tiny\mbox{cal}}}}\circ({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}}}\right]N}}\right)$
$\text{ on $D(0,1)$},$
@@
Since $w$ is uniformly bounded, the family $\mathcal{F}=\mathopen{}\mathclose{{\left\{w_{t}:x\in\mathbb{R}\mapsto w(x){\mathds{1}_{x\leqslant t}};t\in\mathbb{R}}}\right\}$ is $\mathcal{U}(0,1)$-Glivenko-Cantelli, and $(K^{(n)})_{n}$ converges uniformly on $[0,1]$ in probability to the function $K$ given by $K(t)\coloneqq\int_{0}^{t}w(x)\operatorname{d}x$ (Shorack and Wellner, 1986). Then, by continuity at $K$ of the map of Lemma E.8, we get that $(K^{(n)}/K^{(n)}(1))_{n}$ converges uniformly (in probability) to $F^{w}_{{\tiny\mbox{cal}}}$}. Since $\|{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}-K^{(n)}/K^{(n)}(1)\|_{\infty}\leqslant w(+\infty)/((n+1)K^{(n)}(1)+w(+\infty))$ tends uniformly to 0 a.s., $({\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)})_{n}$ converges uniformly (in probability) to $F^{w}_{{\tiny\mbox{cal}}}$}. By continuity of the inverse map at $F^{w}_{{\tiny\mbox{cal}}}$} (see Lemma E.6), we obtain that $({({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}})_{n}$ converges in probability to $({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}$ on $\ell^{\infty}(0,1)$. This proves the first statement.
Next, we turn to prove the second statement. Since $w$ is uniformly bounded, the family $\mathcal{F}=\mathopen{}\mathclose{{\left\{w_{t}:x\in\mathbb{R}\mapsto w(x){\mathds{1}_{x\leqslant t}};t\in\mathbb{R}}}\right\}$ is $\mathcal{U}(0,1)$-Donsker (Shorack and Wellner, 1986). hence there exist $\mathbb{K}=(\mathbb{K}(t))_{t\in[0,1]}$ a Gaussian process such that
$\sqrt{n}\mathopen{}\mathclose{{\left(K^{(n)}-K}}\right)\overset{\mathcal{L}}{\rightarrow}\mathbb{K}\ \text{in $\ell^{\infty}[0,1]$},$
@@
$\mbox{Cov}\mathopen{}\mathclose{{\left(\mathbb{K}(s),\mathbb{K}(t)}}\right)$ $=\int_{0}^{s\wedge t}w^{2}(x)\operatorname{d}x-\int_{0}^{s}w(x)\operatorname{d}x\int_{0}^{t}w(x)\operatorname{d}x;$
$=K(1)^{2}\mathopen{}\mathclose{{\left({\rho^{w}}^{2}{{{V}^{w}_{{\tiny\mbox{cal}}}}}(s\wedge t)-{{F^{w}_{{\tiny\mbox{cal}}}}}(t){{F^{w}_{{\tiny\mbox{cal}}}}}(s)}}\right),$
@@
$\sqrt{n}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left(\frac{K^{(n)}}{K^{(n)}(1)}}}\right)-{{F^{w}_{{\tiny\mbox{cal}}}}}}}\right)\overset{\mathcal{L}}{\rightarrow}\frac{\mathbb{K}}{K(1)}-\frac{\mathbb{K}(1)}{K(1)}{{F^{w}_{{\tiny\mbox{cal}}}}},\ \text{in $\mathcal{C}[0,1]$}.$
@@
$\sqrt{n}\mathopen{}\mathclose{{\left({\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}-{{F^{w}_{{\tiny\mbox{cal}}}}}}}\right)\overset{\mathcal{L}}{\rightarrow}\frac{\mathbb{K}}{K(1)}-\frac{\mathbb{K}(1)}{K(1)}{{F^{w}_{{\tiny\mbox{cal}}}}},\ \text{in $\mathcal{C}[0,1]$}.$
@@
$\sqrt{n}\mathopen{}\mathclose{{\left({\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)-1}-({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}}}\right)\overset{\mathcal{L}}{\rightarrow}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1})^{\prime}\mathopen{}\mathclose{{\left(-\frac{\mathbb{K}\circ({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}}{K(1)}+\frac{\mathbb{K}(1)}{K(1)}I}}\right),\ \text{in $D(0,1)$}.$
To conclude, we identify the covariance of the Gaussian limiting process. For, $1>t\geqslant s>0$,
$\mbox{Cov}\mathopen{}\mathclose{{\left(-\mathbb{K}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(t))/K(1)+\mathbb{K}(1)/K(1)t,-\mathbb{K}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(s))/K(1)+\mathbb{K}(1)/K(1)s}}\right)$
@@
$={\rho^{w}}^{2}\mathopen{}\mathclose{{\left({{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(s))-s{{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(t))-t{{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(s))+ts}}\right).$
@@
$+\mathopen{}\mathclose{{\left[t-{{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(t))}}\right]\mathopen{}\mathclose{{\left[s-{{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(t))}}\right]\Big)$
$=$ ${\rho^{w}}^{2}\mathopen{}\mathclose{{\left({{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(s))\mathopen{}\mathclose{{\left[1-{{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(t))}}\right]+\mathopen{}\mathclose{{\left[t-{{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(t))}}\right]\mathopen{}\mathclose{{\left[s-{{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(t))}}\right]}}\right).$
The first term coincides with the covariance term of $\mathbb{V}\circ{{V}^{w}_{{\tiny\mbox{cal}}}}\circ({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}$ for a standard Brownian bridge $\mathbb{V}$, while the second term is the covariance of the process $\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left[t-{{V}^{w}_{{\tiny\mbox{cal}}}}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(t))}}\right]N}}\right)_{t\in(0,1)}$ with $N\sim\mathcal{N}(0,1)$. This concludes the proof for the second statement.